# Holistic Order Prediction in Natural Scenes

### v1

@@ 1 Introduction
Figure 2: Overview of our holistic approach. We compare the holistic approach to other inference strategies (fig.˜2(a)) and highlight the input-output format discrepancies (fig.˜2(b)).
Few existing approaches explicitly tackle this geometrical scene graph problem Zhan et al. (2020); Zhu et al. (2017); Lee and Park (2022). Yet, they suffer from expensive inference costs and impractical input formats. Specifically, these methods formulate the occlusion and depth ordering prediction tasks a series of instance-wise edge prediction, meaning they require a quadratic amount of passes to obtain the full graph (fig.˜2(a)). Moreover, all the aforementioned arts require the binary masks of all the instances as input (fig.˜2(b)).
@@ 2.1 Order prediction
On the other hand, we reformulate the pairwise edge-level ordering prediction problem to a single adjacency matrix-level prediction. This holistic property of our InstaFormer model yields the full geometrical ordering relations in a unique forward pass, solely from an RGB image given as input (fig.˜2(a)), effectively alleviating the input-output constraint induced by prior approaches while enabling single-pass inference.
@@ 3.2 Preliminaries: pairwise networks
Existing arts Lee and Park (2022); Zhan et al. (2020); Zhu et al. (2017) formulate the occlusion prediction outputs as a binary prediction problem, i.e., $A\rightarrowBB$, $B\rightarrowAA$}. This leads to no occlusion in the simplest case and a bidirectional occlusion in the most complex case. The depth ordering prediction task is formulated as a 3-classes classification problem $A\rightarrowBB$, $B\rightarrowAA$, $A\longleftrightarrowBB$}, with an extra prediction explicitly indicating if $A$ and $B$ share overlapping depth ranges Lee and Park (2022) (fig.˜2(a), pairwise networks).
All these approaches leverage input binary masks, making them cumbersome, especially when ground truth masks are unavailable at test time (fig.˜2(a)).
@@ 3.3 InstaFormer for holistic order prediction
Here, we introduce our main contribution: the InstaFormer family of networks. InstaFormer is capable of holistic order prediction. Given an RGB image as input, it predicts the full order matrices in a single forward pass thanks to the holistic objective formulation and its internal attention mechanism use. The main idea of our method is to create interaction between object queries and latent mask descriptors representing the same objects while carrying complementary information. A dot product, akin to a similarity operator, computes the final geometrical output from these two representations. We invite the reader to follow equations along using fig.˜3.
@@
We want to obtain a global descriptor from each $\mathbf{p}_{i}\in\mathbf{P}$, with $i\in\{0,\dots,n\}$. We found that directly pooling from $\mathbf{P}$ does not provide a comprehensive enough description (table˜2(b)). Thus, we first forward $\mathbf{P}$ to a single transformer layer Dosovitskiy et al. (2021) and max-pool this updated representation instead. To obtain accurate segment-level mask descriptors, we restrict the attention of each $\mathbf{p}_{i}$ to itself using Masked Self-Attention (MSA):
@@
To allow the network to fully benefit from the segmentation latent space of the mask extractor, we attach adapters to every feed-forward network (FFN) in the transformer decoder of the backbone Chen et al. (2022); Kim et al. (2024); Houlsby et al. (2019). We refer the reader to appendix˜E for more information.
@@
4 Experiments
4.1 Baselines
VLM.
We introduce a VLM baseline by converting InstaOrder to InstaOrder-Vqa and prompt the model with occlusion and depth order questions. We use LLaVA Liu et al. (2024a) as it is open-source, popular, and widely used. We run zero-shot evaluation but also fine-tune on our data (denoted with ‡ in table˜1). We refer the reader to appendix˜B and C for more information.
Non-parametric methods.
Foundation models.
Pairwise paradigm.
Mask2Order.
Since pairwise networks cannot perform inference unless a pair of segmentation masks is provided a priori Lee and Park (2022); Zhu et al. (2017); Zhan et al. (2020) (fig.˜2(a), pairwise networks), we introduce Mask2Order to solve this issue in the simplest possible way. It results from the concatenation of a segmentation backbone and a pairwise ordering prediction network (fig.˜2(a), Mask2Order). Motivated by recent advances in this domain Carion et al. (2020); Zhu et al. (2021); Cheng et al. (2021), we select Mask2Former Cheng et al. (2022) as our mask generator. While the Mask2Order framework can obtain geometrical orderings from an RGB image without additional binary masks from the user, this network family still requires multiple forward passes to get the relations between all the instances. We also train InstaOrderNets on top of Mask2Former (denoted with ‡ in table˜1). In this setting, we use the same recipe as Lee and Park (2022) but modernize optimization by switching the optimizer for AwamW and using a cosine schedule.
4.2 Datasets
We run experiments on InstaOrder Lee and Park (2022). We convert InstaOrder to a VQA version, i.e., InstaOrder-Vqa, to evaluate LLaVA Liu et al. (2024b) on occlusion and depth order prediction in zero-shot and finetuned manners (denoted with ‡ in table˜1). We release the conversion script and the dataset with this work. We refer the reader to appendix˜B for more information.
4.3 Evaluation protocol
Metrics.
For occlusion order prediction, we use precision, recall, and F1 scores (higher is better). For depth order prediction, we use the weighted human disagreement rate (WHDR) as proposed in earlier works Lee and Park (2022); Bell et al. (2014) (lower is better). We refer the reader to appendix˜D for detailed definitions.
@@ 4.4 Evaluating naïve baselines
We use the center of the masks as the Y-score and rank them accordingly to obtain the depth orders (table˜1, Y-axis). For the area method, we rank the masks by their size, the bigger the closer (table˜1, Area). These simple heuristics already provide a fair approximation of the geometric order of the elements in the scene.
@@
Zero-shot prompting LLaVA Liu et al. (2024a) scores the least out of all the benchmarked methods (table˜1, LLaVA). We hypothesize that this is because the pre-training of its visual encoder does not enforce precise spatial understanding Radford et al. (2021). We observe notable improvements under the finetuned setting (denoted with ‡), yet, performances remain behind visual experts.
@@
While being incapable of performing occlusion order prediction, we find foundation models to obtain modest results: their performance still lag behind those of specialized networks even the ones that use the same format (table˜1).
@@
Table˜1 confirms that Mask2Order performs worse than InstaOrderNet since it does not benefit from fine-grained segmentation input masks, even when InstaOrderNets are trained on top of Mask2Former (denoted with ‡). This implies that predicting geometrical orderings from self-generated binary masks is a non-trivial task. We further ablate the backbone size of Mask2Order in table˜4(a) and table˜4(b) (appendix˜F). We observe a positive correlation between the size of the backbone and the accuracy results, which seem to confirm our hypothesis stating that pairwise networks are sensitive to the quality of the input mask.
@@ 4.5 Evaluating InstaFormer
Occlusion order recovery.
We evaluate InstaFormero in table˜1. It achieves the best F1 score of the benchmark, even overcoming InstaOrderNeto and InstaOrderNeto,d, which were both trained using a simpler objective and on fine-grained ground truth binary masks. We ablate the segmentation backbone size of our InstaFormero, while keeping the geometric predictor untouched (cf. table˜4(a) (appendix˜F). The more parameters, the better the results. This seems to imply that the ordering prediction results of the geometric predictor depend on the embedding space of the segmentation backbone. In all settings, InstaFormero achieves notable results in comparison to Mask2Ordero. While holistic order prediction is a challenging task, methods can be developed to outperform specialized pairwise models while alleviating the need for input binary masks and reducing the number of forward passes to 1. Most impressively, InstaFormero slightly outperforms InstaOrderNeto,d, the strongest pairwise network using fine-grained ground truth masks as inputs.
Depth order recovery.
We present evaluation results of InstaFormerd in table˜1. InstaFormerd substantially outperforms all baselines on all WHDR (distinct, overlap, all). Similarly to InstaFormero, we also conduct an ablation study related to the backbone size of the segmentation network, while we keep the size of the geometric predictor fixed (table˜4(b), appendix˜F). Surprisingly, the backbone choice has little impact on the final results, which plateaus around the 13.70 mark. Contrary to occlusion order prediction, we hypothesize that depth order prediction does not require edge information and can still operate properly on rougher masks. Nonetheless, InstaFormerd exhibits remarkable results compared to InstaOrderNetd and Mask2Orderd, respectively being on par and beating them even though it operates in a much more challenging setting (i.e. holistic framework: without input masks and by predicting the full depth matrix at once).
@@
We report occlusion order and depth order evaluation for InstaFormero,d in table˜1. InstaFormero,d is competitive with all baselines and sometimes surpasses them, although not using fine-grained ground truth masks as input. It effectively ranks as the strongest depth order prediction network. Contrary to the Mask2Order family, joint occlusion and depth ordering prediction on the InstaFormer family performs similarly to order-specific networks. InstaFormero marginally outperforms InstaFormero,d (table˜1). We suspect the optimization process to be harder for joint occlusion and depth order prediction. In both cases, InstaFormero,d obtains competitive results with InstaOrderNets, especially on depth prediction, even though it was not trained on fine-grained ground truth masks. Interestingly, slightly differently from what was reported in Lee and Park (2022), we observe a substantial improvement in depth performances but not occlusion when performing joint occlusion and depth prediction. Thus, we still conclude that there are benefits to using a joint training scheme both from the computational efficiency and performance standpoints.
Table 2: Ablation study. We report the occlusion and depth order performances on InstaOrder using InstaFormero,d Swin-T.
@@
We benchmark the inference cost of our approach compared to InstaOrderNets, representative of pairwise approaches, in fig.˜5. InstaFormer compares favorably, as its runtime appears constant and memory linear with respect to the number of instances in the image. Details can be found in appendix˜G.
@@ 4.6 Ablation studies
The main idea behind our method is to generate instance-wise interactions between latent object representations. Naïve interaction from object queries to themselves results in low occlusion accuracy and poor depth ordering performances (table˜2(a)). When latent mask descriptors interact with themselves, results for occlusion orders rise, but not for depth orders. Best performances are obtained when combining information from both object queries and latent mask descriptors, as they carry complementary information while referring to the same objects.
@@
Naïvely constructing the latent mask descriptor does not enhance the performances as much as it could in table˜2(b). In fact, directly max-pooling the latent masks does not yield good performance in the depth order prediction task. Instead, simply adding a transformer layer before pooling produces a more comprehensive descriptor, enabling competitive results for occlusion and depth order prediction. As the performances saturate with a single layer, we do not scale more.
@@ 4.7 Qualitative results
InstaFormer can accurately determine geometrical orderings in complex scenes, even for small objects (donuts, cars in fig.˜4) or under in-the-wild inference (K-pop in  fig.˜4, table tennis in fig.˜1). Our network still predicts reasonable relations even under sparse clue inputs (K-Pop example). After text conversion, InstaFormer’s outputs can geometrically ground VLMs and enhance their reasoning in zero-shot prompting (fig.˜1, right). We highlight more results and failure cases in appendix˜H.
@@ Acknowledgments
This work was supported by the IITP grant (RS-2021-II211343: AI Graduate School Program at Seoul National University (5%), RS-2024-00509257: Global AI Frontier Lab (40%), and RS-2025-25442338: AI Star Fellowship Support Program (55%)) funded by the Korea government (MSIT).
@@ Appendix B About InstaOrder-Vqa
To evaluate a wide range of models, we propose a baseline for occlusion and depth order prediction based on LLaVA Liu et al. (2024b, a) (see table˜1 in the main paper). In order to evaluate LLaVA on InstaOrder, we first need to convert the occlusion and depth matrices into a visual question answering (VQA) format.
@@ B.2 Instance-wise disambiguation
However, there still lies ambiguity when two instances in the scene share the same class since then $A_{\text{cls}}=B_{\text{cls}}$. Thus, it is unclear for the model to know which string refers to which instance in the image. To mitigate that issue, we leverage the fact that LLaVA can process bounding box coordinates and input the coordinates of the instances to clarify the position of the object we refer to (fig.˜6).

### v2

@@ 1 Introduction
Figure 2: Overview of our holistic approach. We compare the holistic approach to other inference strategies (Fig. 2(a)) and highlight the input-output format discrepancies (Tab. 2(b)).
Few existing approaches explicitly tackle this geometrical scene graph problem Zhan et al. (2020); Zhu et al. (2017); Lee and Park (2022). Yet, they suffer from expensive inference costs and impractical input formats. Specifically, these methods formulate the occlusion and depth ordering prediction tasks a series of instance-wise edge prediction, meaning they require a quadratic amount of passes to obtain the full graph (Fig. 2(a)). Moreover, all the aforementioned arts require the binary masks of all the instances as input (Fig. 2(b)).
@@ 2.1 Order prediction
On the other hand, we reformulate the pairwise edge-level ordering prediction problem to a single adjacency matrix-level prediction. This holistic property of our InstaFormer model yields the full geometrical ordering relations in a unique forward pass, solely from an RGB image given as input (Fig. 2(a)), effectively alleviating the input-output constraint induced by prior approaches while enabling single-pass inference.
@@ 3.2 Preliminaries: pairwise networks
Existing arts Lee and Park (2022); Zhan et al. (2020); Zhu et al. (2017) formulate the occlusion prediction outputs as a binary prediction problem, i.e., $A\rightarrowBB$, $B\rightarrowAA$}. This leads to no occlusion in the simplest case and a bidirectional occlusion in the most complex case. The depth ordering prediction task is formulated as a 3-classes classification problem $A\rightarrowBB$, $B\rightarrowAA$, $A\longleftrightarrowBB$}, with an extra prediction explicitly indicating if $A$ and $B$ share overlapping depth ranges Lee and Park (2022) (Fig. 2(a), pairwise networks).
All these approaches leverage input binary masks, making them cumbersome, especially when ground truth masks are unavailable at test time (Fig. 2(a)).
@@ 3.3 InstaFormer for holistic order prediction
Here, we introduce our main contribution: the InstaFormer family of networks. InstaFormer is capable of holistic order prediction. Given an RGB image as input, it predicts the full order matrices in a single forward pass thanks to the holistic objective formulation and its internal attention mechanism use. The main idea of our method is to create interaction between object queries and latent mask descriptors representing the same objects while carrying complementary information. A dot product, akin to a similarity operator, computes the final geometrical output from these two representations. We invite the reader to follow equations along using Fig. 3.
@@
We want to obtain a global descriptor from each $\mathbf{p}_{i}\in\mathbf{P}$, with $i\in\{0,\dots,n\}$. We found that directly pooling from $\mathbf{P}$ does not provide a comprehensive enough description (Tab. 2(b)). Thus, we first forward $\mathbf{P}$ to a single transformer layer Dosovitskiy et al. (2021) and max-pool this updated representation instead. To obtain accurate segment-level mask descriptors, we restrict the attention of each $\mathbf{p}_{i}$ to itself using Masked Self-Attention (MSA):
@@
To allow the network to fully benefit from the segmentation latent space of the mask extractor, we attach adapters to every feed-forward network (FFN) in the transformer decoder of the backbone Chen et al. (2022); Kim et al. (2024); Houlsby et al. (2019). We refer the reader to Appendix˜D for more information.
@@
4 Experiments
4.1 Baselines
VLM.
We introduce a VLM baseline by converting InstaOrder to InstaOrder-Vqa and prompt the model with occlusion and depth order questions. We use LLaVA Liu et al. (2024a) as it is open-source, popular, and widely used. We run zero-shot evaluation but also fine-tune on our data (denoted with ‡ in Tab. 1). We refer the reader to Appendix˜B and C for more information.
Non-parametric methods.
Foundation models.
Pairwise paradigm.
Mask2Order.
Since pairwise networks cannot perform inference unless a pair of segmentation masks is provided a priori Lee and Park (2022); Zhu et al. (2017); Zhan et al. (2020) (Fig. 2(a), pairwise networks), we introduce Mask2Order to solve this issue in the simplest possible way. It results from the concatenation of a segmentation backbone and a pairwise ordering prediction network (Fig. 2(a), Mask2Order). Motivated by recent advances in this domain Carion et al. (2020); Zhu et al. (2021); Cheng et al. (2021), we select Mask2Former Cheng et al. (2022) as our mask generator. While the Mask2Order framework can obtain geometrical orderings from an RGB image without additional binary masks from the user, this network family still requires multiple forward passes to get the relations between all the instances. We also train InstaOrderNets on top of Mask2Former (denoted with ‡ in Tab. 1). In this setting, we use the same recipe as Lee and Park (2022) but modernize optimization by switching the optimizer for AwamW and using a cosine schedule.
4.2 Datasets
We run experiments on InstaOrder Lee and Park (2022). We convert InstaOrder to a VQA version, i.e., InstaOrder-Vqa, to evaluate LLaVA Liu et al. (2024b) on occlusion and depth order prediction in zero-shot and finetuned manners (denoted with ‡ in Tab. 1). We release the conversion script and the dataset with this work. We refer the reader to Appendix˜B for more information.
4.3 Evaluation protocol
Metrics.
Following previous works Zhan et al. (2020); Zhu et al. (2017); Lee and Park (2022), we evaluate the performance of our networks on the task of occlusion ordering prediction by reporting precision, recall, and F1-score. We compare each element of the predicted ordering matrix to each element of the ground truth matrix. Given two instances $A$ and $B$, the aforementioned metrics are expressed as follows:
where $w_{AB}=\frac{2}{\text{count}_{AB}}$.
@@ 4.4 Evaluating naïve baselines
We use the center of the masks as the Y-score and rank them accordingly to obtain the depth orders (Tab. 1, Y-axis). For the area method, we rank the masks by their size, the bigger the closer (Tab. 1, Area). These simple heuristics already provide a fair approximation of the geometric order of the elements in the scene.
@@
Zero-shot prompting LLaVA Liu et al. (2024a) scores the least out of all the benchmarked methods (Tab. 1, LLaVA). We hypothesize that this is because the pre-training of its visual encoder does not enforce precise spatial understanding Radford et al. (2021). We observe notable improvements under the finetuned setting (denoted with ‡), yet, performances remain behind visual experts.
@@
While being incapable of performing occlusion order prediction, we find foundation models to obtain modest results: their performance still lag behind those of specialized networks even the ones that use the same format (Tab. 1).
@@
Table˜1 confirms that Mask2Order performs worse than InstaOrderNet since it does not benefit from fine-grained segmentation input masks, even when InstaOrderNets are trained on top of Mask2Former (denoted with ‡). This implies that predicting geometrical orderings from self-generated binary masks is a non-trivial task. We further ablate the backbone size of Mask2Order in Tab. 4(a) and 4(b) (Appendix˜E). We observe a positive correlation between the size of the backbone and the accuracy results, which seem to confirm our hypothesis stating that pairwise networks are sensitive to the quality of the input mask.
@@ 4.5 Evaluating InstaFormer
Occlusion order recovery.
We evaluate InstaFormero in Tab. 1. It achieves the best F1 score of the benchmark, even overcoming InstaOrderNeto and InstaOrderNeto,d, which were both trained using a simpler objective and on fine-grained ground truth binary masks. We ablate the segmentation backbone size of our InstaFormero, while keeping the geometric predictor untouched (cf. Tab. 4(a) (Appendix˜E). The more parameters, the better the results. This seems to imply that the ordering prediction results of the geometric predictor depend on the embedding space of the segmentation backbone. In all settings, InstaFormero achieves notable results in comparison to Mask2Ordero. While holistic order prediction is a challenging task, methods can be developed to outperform specialized pairwise models while alleviating the need for input binary masks and reducing the number of forward passes to 1. Most impressively, InstaFormero slightly outperforms InstaOrderNeto,d, the strongest pairwise network using fine-grained ground truth masks as inputs.
Depth order recovery.
We present evaluation results of InstaFormerd in Tab. 1. InstaFormerd substantially outperforms all baselines on all WHDR (distinct, overlap, all). Similarly to InstaFormero, we also conduct an ablation study related to the backbone size of the segmentation network, while we keep the size of the geometric predictor fixed (Tab. 4(b), Appendix˜E). Surprisingly, the backbone choice has little impact on the final results, which plateaus around the 13.70 mark. Contrary to occlusion order prediction, we hypothesize that depth order prediction does not require edge information and can still operate properly on rougher masks. Nonetheless, InstaFormerd exhibits remarkable results compared to InstaOrderNetd and Mask2Orderd, respectively being on par and beating them even though it operates in a much more challenging setting (i.e. holistic framework: without input masks and by predicting the full depth matrix at once).
@@
We report occlusion order and depth order evaluation for InstaFormero,d in Tab. 1. InstaFormero,d is competitive with all baselines and sometimes surpasses them, although not using fine-grained ground truth masks as input. It effectively ranks as the strongest depth order prediction network. Contrary to the Mask2Order family, joint occlusion and depth ordering prediction on the InstaFormer family performs similarly to order-specific networks. InstaFormero marginally outperforms InstaFormero,d (Tab. 1). We suspect the optimization process to be harder for joint occlusion and depth order prediction. In both cases, InstaFormero,d obtains competitive results with InstaOrderNets, especially on depth prediction, even though it was not trained on fine-grained ground truth masks. Interestingly, slightly differently from what was reported in Lee and Park (2022), we observe a substantial improvement in depth performances but not occlusion when performing joint occlusion and depth prediction. Thus, we still conclude that there are benefits to using a joint training scheme both from the computational efficiency and performance standpoints.
Table 2: Ablation study. Occlusion and depth order performances on InstaOrder using InstaFormero,d Swin-T.
@@
We benchmark the inference cost of our approach compared to InstaOrderNets, representative of pairwise approaches, in Fig. 5. InstaFormer compares favorably, as its runtime appears constant and memory linear with respect to the number of instances in the image. Details can be found in Appendix˜F.
@@ 4.6 Ablation studies
The main idea behind our method is to generate instance-wise interactions between latent object representations. Naïve interaction from object queries to themselves results in low occlusion accuracy and poor depth ordering performances (Tab. 2(a)). When latent mask descriptors interact with themselves, results for occlusion orders rise, but not for depth orders. Best performances are obtained when combining information from both object queries and latent mask descriptors, as they carry complementary information while referring to the same objects.
@@
Naïvely constructing the latent mask descriptor does not enhance the performances as much as it could in Tab. 2(b). In fact, directly max-pooling the latent masks does not yield good performance in the depth order prediction task. Instead, simply adding a transformer layer before pooling produces a more comprehensive descriptor, enabling competitive results for occlusion and depth order prediction. As the performances saturate with a single layer, we do not scale more.
@@ 4.7 Qualitative results
InstaFormer can accurately determine geometrical orderings in complex scenes, even for small objects (donuts, cars in Fig. 4) or under in-the-wild inference (K-pop in Fig. 4, table tennis in Fig. 1). Our network still predicts reasonable relations even under sparse clue inputs (K-Pop example). After text conversion, InstaFormer’s outputs can geometrically ground VLMs and enhance their reasoning in zero-shot prompting (Fig. 1, right). We provide diverse qualitative results of our model on InstaOrder in Fig. 10, in Appendix˜G. We observe that our model predicts sane orderings in various scene types. Ground truths for all the predictions are available in Fig. 9 and Fig. 11 in the Appendices.
(a) Instance-mixup.
(b) Segmentation failure.
Figure 6: Failure cases of InstaFormero,d. We highlight two failure cases of our network: the “instance-mixup” (Fig. 6(a)) and the “segmentation failure” (Fig. 6(b)).
4.8 Failure cases
We highlight two cases in which InstaFormer struggles to predict accurate relationships between instances Fig. 6. We name these situation “instance mixup” and “segmentation failure”. In the “instance-mixup” case, the “Person” instances related to the “Ski” are shared instead of being targeted solely on “Person1”. In the “segmentation failure” case, a singular instance “Elephant” is wrongfully segmented multiple times, leading to multiple false positive nodes in the geometrical graphs.
@@ Acknowledgments
This work was supported by the IITP grant (RS-2021-II211343: AI Graduate School Program at Seoul National University (5%), RS-2024-00509257: Global AI Frontier Lab (40%), and RS-2025-25442338: AI Star Fellowship Support Program (55%)) funded by the Korean government (MSIT).
@@ Appendix B About InstaOrder-Vqa
To evaluate a wide range of models, we propose a baseline for occlusion and depth order prediction based on LLaVA Liu et al. (2024b, a) (see Tab. 1 in the main paper). In order to evaluate LLaVA on InstaOrder, we first need to convert the occlusion and depth matrices into a visual question answering (VQA) format.
@@ B.2 Instance-wise disambiguation
However, there still lies ambiguity when two instances in the scene share the same class since then $A_{\text{cls}}=B_{\text{cls}}$. Thus, it is unclear for the model to know which string refers to which instance in the image. To mitigate that issue, we leverage the fact that LLaVA can process bounding box coordinates and input the coordinates of the instances to clarify the position of the object we refer to (Fig. 7).